# Linearization in magnetoelasticity

### v1

@@ 2. Auxiliary results
Define a homogeneous Sobolev space $L^{1,2}(\mathbb{R}^{d})$ as the space of $\varphi\in L^{2}_{loc}(\mathbb{R}^{d})$ such that $\nabla\varphi\in L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Then magnetic potential $v_{m}\in L^{1,2}(\mathbb{R}^{d})$ generated by magnetization $m$ exists, unique up to a constant, and continuously depends on $m$.
@@
Moreover, if $f_{j}\to f$ strongly in $L^{2}$, then
$\|\nabla v_{f_{j}}\|_{L^{2}}\to\|\nabla v_{f}\|_{L^{2}}.$
@@
This proposition is classical, see for example [2, Proposition 8.8 & Theorem 8.9] or [23, Proof of Theorem 4.1]. ∎
@@
    $|\nabla u|_{O}<1\Rightarrow\det\nabla y>0.$
@@
To establish compactness, we need the following geometric rigidity, proven in [15, Theorem 3.1] for $p=2$ and in [9, Section 2.4] for $p\in(1,+\infty)$.
@@
The particular choice of $R$ as in (2.2) can be ensured arguing, e.g., as in [1, Theorem 3.1]. ∎
@@ 3. Compactness
Under the assumptions of Theorem 1.2, let $(u_{\varepsilon},m_{\varepsilon})\in\mathcal{A}^{w}_{\varepsilon}$ be such that, for $y_{\varepsilon}:=id+\varepsilon u_{\varepsilon}$, it holds
@@
Then, there exist $(u,m)\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})\times W^{1,2}(\Omega;\mathbb{S}^{d-1})$ such that, up to a subsequence,
@@
Under the assumptions of Theorem 1.1, there exists a positive constant $C=C(\Omega,d,p)>0$ such that for every $\varepsilon\in(0,1)$ and every $(u,m)\in\mathcal{A}^{w}_{\varepsilon}$ it holds
(3.2) $\varepsilon^{2}\mathcal{G}_{\varepsilon}(u,m)\geq C\int_{\Omega}g_{p}(\operatorname{dist}(\nabla y_{\varepsilon},SO(d))\,{\,\rm d}x-C\varepsilon^{2},$
@@
(3.5) $\int_{\Omega}|\nabla u|^{2}{\,\rm d}x\leq C\Bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Gamma}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
(3.6) $\int_{\Omega}|\varepsilon\nabla u|^{p}{\,\rm d}x\leq C\varepsilon^{2}\bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Gamma}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)}.$
@@
Arguing as in [10, Proposition 3.4], we get that
(3.8) $|R_{\varepsilon}-I|^{2}\leq C\varepsilon^{2}\bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Gamma}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
@@
$\leq C\bigg{(}\varepsilon^{2}\mathcal{G}_{\varepsilon}(u,m)+\varepsilon^{2}+|R_{\varepsilon}-I|^{2}\bigg{)}\leq C\varepsilon^{2}\bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Gamma}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
@@
We are now ready to prove the compactness result of Proposition 3.1.
@@
$\int_{\Omega}\frac{1}{(\det\nabla y_{\varepsilon})^{a}}{\,\rm d}x$ $=\int_{\Omega}\frac{1}{(\det\Lambda_{\varepsilon})^{a}}{\,\rm d}x$
@@
Furthermore, we recall that (3.6) implies that
(3.9) $\|\nabla y_{\varepsilon}-I\|_{L^{p}}\leq C\varepsilon^{\frac{2}{p}}$
for some positive constant $C$ independent of $\varepsilon$.
@@
Without loss of generality, we may assume that $\Omega_{\delta}$ is a Lipschitz set. Since $y_{\varepsilon}\to id$ in $W^{1,p}(\Omega;\mathbb{R}^{d})$ and $p>d$, we have that $y_{\varepsilon}\to id$ in $C^{0}(\overline{\Omega};\mathbb{R}^{d})$. Hence, for $\varepsilon>0$ small enough, we may assume that $\Omega_{\delta}\subseteq y_{\varepsilon}(\Omega)$. By change of variables, using the equality $|m_{\varepsilon}\circ y_{\varepsilon}|\det\nabla y_{\varepsilon}=1$ a.e. in $\Omega$, we get that
@@
Therefore, $\mu_{\varepsilon}^{\delta}$ is bounded in $W^{1,2}(\Omega_{\delta};\mathbb{R}^{d})$ uniformly for all $\delta>0$. By a diagonal argument, we find $m\in W^{1,2}_{loc}(\Omega;\mathbb{R}^{d})$ such that $\mu_{\varepsilon}^{\delta_{\varepsilon}}\rightharpoonup m$ weakly in $W^{1,2}_{loc}(\Omega;\mathbb{R}^{d})$. While the Sobolev Embedding Theorems provide that
(3.12) $m_{\varepsilon}(x)=\mu_{\varepsilon}^{\delta}(x)\to m(x)\qquad\text{for a.e.~{}}x\in\Omega_{\delta}.$
@@
Passing to the liminf by $\varepsilon$ and then to the limit by $\delta$ in (3.10) and (3.11), we further deduce that $m\in W^{1,2}(\Omega;\mathbb{R}^{d})$. Note further that the second line in (3.10) ensures that $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}$ is bounded in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and thus converges weakly to some $\mu\in L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Using pointwise convergence (3.12), we conclude $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}(x)\to\chi_{\Omega}m(x)$ a.e. in $\mathbb{R}^{d}$. Therefore $\mu=\chi_{\Omega}m$ a.e., and, up to a subsequence,
(3.13) $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\rightharpoonup\chi_{\Omega}m\quad\text{ weakly in }L^{2}(\mathbb{R}^{d};\mathbb{R}^{d}).$
@@
Further, applying the Lebesgue–Vitali Convergence Theorem (see [3, Th. 4.5.4] or [22, Ch. 6, Ex. 10]) to $|\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}|^{2}$, we obtain convergence of norms
$\|\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\|_{L^{2}}\to\|\chi_{\Omega}m\|_{L^{2}}.$
@@
This, together with weak convergence (3.13) guarantees strong convergence $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m$ in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Since the weak convergence in $L^{2}(\Omega;\mathbb{R}^{d})$ of $m_{\varepsilon}\chi_{y_{\varepsilon}(\Omega)}$ to $m$ has been already shown, we deduce the strong $L^{2}$-convergence.
@@
By (3.1) we know that $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}$ is bounded in $L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$. For every test $\varphi\in L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$ we consider
$\int_{\mathbb{R}^{d}}$ $(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m):\varphi{\,\rm d}x$
$=\int_{\Omega_{\delta}}(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m):\varphi{\,\rm d}x+\int_{\mathbb{R}^{d}\setminus\Omega_{\delta}}(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m):\varphi{\,\rm d}x.$
@@
Then, the first integral tends to 00 since $\Omega_{\delta}\subseteq y_{\varepsilon}(\Omega)\cap\Omega$, $m_{\varepsilon}=\mu_{\varepsilon}^{\delta}$ in $\Omega_{\delta}$, and $\mu_{\varepsilon}^{\delta}\rightharpoonup m$ weakly in $W^{1,2}(\Omega_{\delta};\mathbb{R}^{d})$. As for the second integral, we have that
@@
$\bigg{|}\int_{\mathbb{R}^{d}\setminus\Omega_{\delta}}(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m):\varphi{\,\rm d}x\bigg{|}\leq C\|\varphi\|_{L^{2}(C_{\delta})},$
@@
where $C_{\delta}=\{x\in\mathbb{R}^{d}:\,\operatorname{dist}(x,\partial\Omega)<\delta\}$. Since $\mathcal{L}^{d}(C_{\delta})\to 0$ as $\delta\to 0$, we infer that $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}$ converges to $\chi_{\Omega}\nabla m$ weakly in $L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$.
@@
We now show that $m_{\varepsilon}\circ y_{\varepsilon}$ converges to $m$ in $L^{1}(\Omega;\mathbb{R}^{d})$. For $\delta>0$ and $\varepsilon\ll 1$ we write
(3.14) $\int_{\Omega_{\delta}}|m_{\varepsilon}\circ y_{\varepsilon}-m|{\,\rm d}x\leq\int_{\Omega_{\delta}}|m_{\varepsilon}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x+\int_{\Omega_{\delta}}|\mu_{\varepsilon}^{\delta}-m|{\,\rm d}x.$
@@
For the second integral on the right-hand side of (3.14) we simply have that $\mu_{\varepsilon}^{\delta}\to m$ in $L^{2}(\Omega_{\delta};\mathbb{R}^{d})$. As for the first term on the right-hand side of (3.14), we notice that $m_{\varepsilon}\circ y_{\varepsilon}$ and $\mu_{\varepsilon}^{\delta}$ are equi-integrable in $\Omega_{\delta}$. Thus, for $\eta>0$ there exists $\rho>0$ such that for every $A\subseteq\Omega_{\delta}$ measurable
@@
$\limsup_{\varepsilon\to 0}\mathcal{L}^{d}(\Omega_{\delta}\setminus A_{\delta,\varepsilon})$ $\leq\limsup_{\varepsilon\to 0}\mathcal{L}^{d}(\Omega_{\overline{\delta}})-\mathcal{L}^{d}(A_{\delta,\varepsilon})$
$=\limsup_{\varepsilon\to 0}\mathcal{L}^{d}(\Omega_{\overline{\delta}}\setminus\Omega_{\delta})-\int_{A_{\delta,\varepsilon}}(\det\nabla y_{\varepsilon}-1){\,\rm d}x<\rho.$
@@
(3.18) $\limsup_{\varepsilon\to 0}\int_{\Omega_{\delta}\cap A_{\delta,\varepsilon}\setminus(X_{\delta,\varepsilon}\cap F_{\delta,\varepsilon})}|m_{\varepsilon}^{\delta}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x<\eta.$
@@
$\limsup_{\varepsilon\to 0}\int_{\Omega_{\delta}}|m_{\varepsilon}\circ y_{\varepsilon}-m|{\,\rm d}x<3\eta.$
Since $\eta>0$ is arbitrary, this implies that $m_{\varepsilon}\circ y_{\varepsilon}\to m$ in $L^{1}(\Omega_{\delta};\mathbb{R}^{d})$ for every $\delta>0$. By the equi-integrability of $m_{\varepsilon}\circ y_{\varepsilon}$, this entails the convergence in $L^{1}(\Omega;\mathbb{R}^{d})$. Furthermore, since strong $L^{1}$-convergence implies convergence a.e., we have $m_{\varepsilon}\circ y_{\varepsilon}\det\nabla y_{\varepsilon}\to m$ a.e. and $|m_{\varepsilon}\circ y_{\varepsilon}|\det\nabla y_{\varepsilon}=|m|=1$ a.e. in $\Omega$. Thus, $m\in W^{1,2}(\Omega;\mathbb{S}^{d-1})$ and $m_{\varepsilon}\circ y_{\varepsilon}\det\nabla y_{\varepsilon}\to m$ in $L^{r}(\Omega;\mathbb{R}^{d})$ for any $r\in[1,+\infty)$.

### v2

@@ 2. Auxiliary results
The homogeneous Sobolev space $L^{1,2}(\mathbb{R}^{d})$ is defined as the space of $\varphi\in L^{2}_{loc}(\mathbb{R}^{d})$ such that $\nabla\varphi\in L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. We equip the space $L^{1,2}(\mathbb{R}^{d})$ with a seminorm $\|\nabla\varphi\|_{L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})}$. It can be as well supplied with a norm $\|\varphi\|_{L^{1,2}(\mathbb{R}^{d})}:=\|\nabla\varphi\|_{L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})}+\|\varphi\|_{L^{2}(\omega)}$, where $\omega\subset\mathbb{R}^{d}$ is arbitrary bounded open nonempty set. For more details, the interested reader is referred to [34, Chapter 1]. The magnetic potential $v_{m}\in L^{1,2}(\mathbb{R}^{d})$ generated by magnetization $m$ exists, unique up to a constant, and continuously depends on $m$.
@@
Moreover, if $f_{j}\to f$ strongly in $L^{2}$, then $\nabla v_{f_{j}}\to\nabla v_{f}$ strongly in $L^{2}(\mathbb{R}^{d},\mathbb{R}^{d})$.
@@
This proposition is classical, see for example [5, Proposition 8.8 & Theorem 8.9] or [39, Proof of Theorem 4.1]. ∎
@@
    $|\nabla u(x)|_{O}<1\Rightarrow\det\nabla y(x)>0.$
@@
To establish compactness, we need the following geometric rigidity, proven in [26, Theorem 3.1] for $p=2$ and in [16, Section 2.4] for $p\in(1,+\infty)$.
@@
The particular choice of $R$ as in (2.2) can be ensured arguing, e.g., as in [4, Theorem 3.1]. ∎
@@ 3. Compactness
Under the assumptions of Theorem 1.3, let $(u_{\varepsilon},m_{\varepsilon})\in\mathcal{A}^{w}_{\varepsilon}$ be such that
@@
Then, there exists $(u_{0},m_{0})\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})\times W^{1,2}(\Omega;\mathbb{S}^{d-1})$ such that, up to a subsequence,
@@
Under the assumptions of Theorem 1.1, there exist positive constants $C_{1}=C_{1}(\Omega,d,p)>0$ and $C_{2}=C_{2}(\Omega,d,p)>0$ such that for every $\varepsilon\in(0,1)$ and every $(u,m)\in\mathcal{A}^{w}_{\varepsilon}$ it holds
(3.2) $\varepsilon^{2}\mathcal{G}_{\varepsilon}(u,m)\geq C_{1}\int_{\Omega}g_{p}(\operatorname{dist}(\nabla y_{\varepsilon},SO(d))\,{\,\rm d}x-C_{2}\varepsilon^{2},$
@@
(3.5) $\int_{\Omega}|\nabla u|^{2}{\,\rm d}x\leq C\Bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\int_{\Gamma}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
(3.6) $\int_{\Omega}|\varepsilon\nabla u|^{p}{\,\rm d}x\leq C\varepsilon^{2}\bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\int_{\Gamma}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)}.$
@@
Arguing as in [17, Proposition 3.4], we get that
(3.8) $|R_{\varepsilon}-I|^{2}\leq C\varepsilon^{2}\bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\int_{\Gamma}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
@@
$\leq C\bigg{(}\varepsilon^{2}\mathcal{G}_{\varepsilon}(u,m)+\varepsilon^{2}+|R_{\varepsilon}-I|^{2}\bigg{)}\leq C\varepsilon^{2}\bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\int_{\Gamma}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
@@
We are now ready to prove the compactness result of Proposition 3.1. We closely follow the arguments of [10, Proposition 4.3] (see also [13, Proposition 3.7]).
@@
Let us notice that the Poincaré inequality and Lemma 3.3 guarantee, up to a subsequence, weak convergence $u_{\varepsilon}\rightharpoonup u_{0}$ in $W^{1,2}(\Omega,\mathbb{R}^{d})$. Moreover, (3.6) implies that
(3.9) $\|\nabla y_{\varepsilon}-I\|_{L^{p}}\leq C\varepsilon^{\frac{2}{p}}$
for some positive constant $C$ independent of $\varepsilon$. And thus, strong convergence $y_{\varepsilon}\to id$ in $W^{1,p}(\Omega,\mathbb{R}^{d})$.
@@
(3.10) $\int_{\Omega}\frac{1}{(\det\nabla y_{\varepsilon})^{a}}{\,\rm d}x$ $=\int_{\Omega}\frac{1}{(\det\Lambda_{\varepsilon})^{a}}{\,\rm d}x$
@@
Without loss of generality, we may assume that $\Omega_{\delta}$ is a Lipschitz set. Since $y_{\varepsilon}\to id$ in $W^{1,p}(\Omega;\mathbb{R}^{d})$ and $p>d$, we have that $y_{\varepsilon}\to id$ in $C^{0}(\overline{\Omega};\mathbb{R}^{d})$. Hence, for $\varepsilon>0$ small enough, we may assume that $\Omega_{\delta}\subseteq y_{\varepsilon}(\Omega)$. Without going into details about the degree theory, see [25], let us provide rather a standard argument to prove this fact. For $y_{\varepsilon}$, let us define the topological degree $\deg(x,y_{\varepsilon},\Omega)$ as a topological degree of its continuous representative. Notice that $h_{t}(x):=ty_{\varepsilon}(x)+(1-t)x$ is a homotopy between $y_{\varepsilon}$ and $id$, and by the uniform convergence $y_{\varepsilon}\to id$, for any $x\in\Omega_{\delta}$ it holds that $x\not\in y_{\varepsilon}(\partial\Omega)$ for any small enough $\varepsilon$. Thus, $\deg(x,y_{\varepsilon},\Omega)=\deg(x,id,\Omega)>0$, in other words, $x\in y_{\varepsilon}(\Omega)$.
Next, by change of variables, using the equality $|m_{\varepsilon}\circ y_{\varepsilon}|\det\nabla y_{\varepsilon}=1$ a.e. in $\Omega$, we get that
@@
Therefore, $\mu_{\varepsilon}^{\delta}$ is bounded in $W^{1,2}(\Omega_{\delta};\mathbb{R}^{d})$ uniformly for all $\delta>0$. By a diagonal argument, we find $m_{0}\in W^{1,2}_{loc}(\Omega;\mathbb{R}^{d})$ such that $\mu_{\varepsilon}^{\delta}\rightharpoonup m_{0}$ weakly in $W^{1,2}_{loc}(\Omega;\mathbb{R}^{d})$. While the Sobolev Embedding Theorems provide that
(3.13) $m_{\varepsilon}(x)=\mu_{\varepsilon}^{\delta}(x)\to m_{0}(x)\qquad\text{for a.e.~{}}x\in\Omega_{\delta}.$
@@
Passing to the liminf as $\varepsilon\to 0$ and then to the limit as $\delta\to 0$ in (3.11) and (3.12), we further deduce that $m_{0}\in W^{1,2}(\Omega;\mathbb{R}^{d})$. Note further that the second line in (3.11) ensures that $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}$ is bounded in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and thus converges weakly to some $\mu\in L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Using pointwise convergence (3.13), we conclude $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}(x)\to\chi_{\Omega}m_{0}(x)$ a.e. in $\mathbb{R}^{d}$. Therefore $\mu=\chi_{\Omega}m_{0}$ a.e., and, up to a subsequence,
(3.14) $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\rightharpoonup\chi_{\Omega}m_{0}\quad\text{ weakly in }L^{2}(\mathbb{R}^{d};\mathbb{R}^{d}).$
@@
Now it is not difficult to see that $|\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}|^{2}$ converges in measure and, moreover, is equi-integrable by (3.10) and (3.11). Thus, applying the Lebesgue–Vitali Convergence Theorem (see [8, Th. 4.5.4] or [38, Ch. 6, Ex. 10]) to $|\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}|^{2}$, we obtain convergence of norms
$\|\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\|_{L^{2}}\to\|\chi_{\Omega}m_{0}\|_{L^{2}}.$
@@
This, together with weak convergence (3.14) guarantees strong convergence $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m_{0}$ in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Since the weak convergence in $L^{2}(\Omega;\mathbb{R}^{d})$ of $m_{\varepsilon}\chi_{y_{\varepsilon}(\Omega)}$ to $m_{0}$ has been already shown, we deduce the strong $L^{2}$-convergence.
@@
By (3.1) we know that $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}$ is bounded in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d\times d})$. For every test $\varphi\in L^{2}(\mathbb{R}^{d};\mathbb{R}^{d\times d})$ we consider
$\int_{\mathbb{R}^{d}}$ $(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m_{0}):\varphi{\,\rm d}x$
$=\int_{\Omega_{\delta}}(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m_{0}):\varphi{\,\rm d}x+\int_{\mathbb{R}^{d}\setminus\Omega_{\delta}}(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m_{0}):\varphi{\,\rm d}x.$
@@
Then, the first integral tends to 00 since $\Omega_{\delta}\subseteq y_{\varepsilon}(\Omega)\cap\Omega$, $m_{\varepsilon}=\mu_{\varepsilon}^{\delta}$ in $\Omega_{\delta}$, and $\mu_{\varepsilon}^{\delta}\rightharpoonup m_{0}$ weakly in $W^{1,2}(\Omega_{\delta};\mathbb{R}^{d})$. As for the second integral, we have that
@@
$\bigg{|}\int_{\mathbb{R}^{d}\setminus\Omega_{\delta}}(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m_{0}):\varphi{\,\rm d}x\bigg{|}\leq C\|\varphi\|_{L^{2}(C_{\delta})},$
@@
where $C_{\delta}=\{x\in\mathbb{R}^{d}:\,\operatorname{dist}(x,\partial\Omega)<\delta\}$. Since $\mathcal{L}^{d}(C_{\delta})\to 0$ as $\delta\to 0$, we infer that $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}$ converges to $\chi_{\Omega}\nabla m_{0}$ weakly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d\times d})$.
@@
We now show that $m_{\varepsilon}\circ y_{\varepsilon}$ converges to $m_{0}$ in $L^{1}(\Omega;\mathbb{R}^{d})$. For $\delta>0$ and $\varepsilon\ll 1$ we write
(3.15) $\int_{\Omega_{\delta}}|m_{\varepsilon}\circ y_{\varepsilon}-m_{0}|{\,\rm d}x\leq\int_{\Omega_{\delta}}|m_{\varepsilon}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x+\int_{\Omega_{\delta}}|\mu_{\varepsilon}^{\delta}-m_{0}|{\,\rm d}x.$
@@
For the second integral on the right-hand side of (3.15) we simply have that $\mu_{\varepsilon}^{\delta}\to m_{0}$ in $L^{2}(\Omega_{\delta};\mathbb{R}^{d})$. As for the first term on the right-hand side of (3.15), we notice that $m_{\varepsilon}\circ y_{\varepsilon}$ and $\mu_{\varepsilon}^{\delta}$ are equi-integrable in $\Omega_{\delta}$. Thus, for $\eta>0$ there exists $\rho>0$ such that for every $A\subseteq\Omega_{\delta}$ measurable
@@
$\limsup_{\varepsilon\to 0}\mathcal{L}^{d}(\Omega_{\delta}\setminus A_{\delta,\varepsilon})$ $\leq\limsup_{\varepsilon\to 0}\Big{(}\mathcal{L}^{d}(\Omega_{\overline{\delta}})-\mathcal{L}^{d}(A_{\delta,\varepsilon})\Big{)}$
$=\limsup_{\varepsilon\to 0}\Bigg{(}\mathcal{L}^{d}(\Omega_{\overline{\delta}}\setminus\Omega_{\delta})-\int_{A_{\delta,\varepsilon}}(\det\nabla y_{\varepsilon}-1){\,\rm d}x\Bigg{)}<\rho.$
@@
(3.19) $\limsup_{\varepsilon\to 0}\int_{\Omega_{\delta}\cap A_{\delta,\varepsilon}\setminus(X_{\delta,\varepsilon}\cap G_{\delta,\varepsilon}})|m_{\varepsilon}^{\delta}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x<\eta.$
@@
$\limsup_{\varepsilon\to 0}\int_{\Omega_{\delta}}|m_{\varepsilon}\circ y_{\varepsilon}-m_{0}|{\,\rm d}x<3\eta.$
Since $\eta>0$ is arbitrary, this implies that $m_{\varepsilon}\circ y_{\varepsilon}\to m_{0}$ in $L^{1}(\Omega_{\delta};\mathbb{R}^{d})$ for every $\delta>0$. By the equi-integrability of $m_{\varepsilon}\circ y_{\varepsilon}$, this entails the convergence in $L^{1}(\Omega;\mathbb{R}^{d})$. Furthermore, since strong $L^{1}$-convergence implies convergence a.e., we have $m_{\varepsilon}\circ y_{\varepsilon}\det\nabla y_{\varepsilon}\to m_{0}$ a.e. and $|m_{\varepsilon}\circ y_{\varepsilon}|\det\nabla y_{\varepsilon}=|m_{0}|=1$ a.e. in $\Omega$. Thus, $m\in W^{1,2}(\Omega;\mathbb{S}^{d-1})$ and $m_{\varepsilon}\circ y_{\varepsilon}\det\nabla y_{\varepsilon}\to m_{0}$ in $L^{r}(\Omega;\mathbb{R}^{d})$ for any $r\in[1,+\infty)$. ∎